OpenStudy (writerforever18):

Last year, Hans had $30,000 to invest. He invested some of it in an account that paid 9% simple interest per year, and he invested the rest in an account that paid 5% simple interest per year. After one year, he received a total of $2060 in interest. How much did he invest in each account?

OpenStudy (malijhaa):

What Do Yhu Think It Is?

OpenStudy (writerforever18):

I know that the total is 2060 dollars after 1 year.

OpenStudy (writerforever18):

And the year before, he 30,000 dollars to invest in. In the first account he had invested 9% and in the second he invested 5%.

OpenStudy (writerforever18):

But my problem is putting the problem together in order to solve it.

OpenStudy (writerforever18):

Could you help me with that please?

OpenStudy (malijhaa):

So Yhu Know Yhu Have To Figure Out The Percentages Right?

OpenStudy (malijhaa):

Yhu Have To Find The Number They Represent Before Yhu Can Do Anything Else

OpenStudy (writerforever18):

So wouldnt that mean that 9% represents account 1. 5% represents account 2. 2060 dollars represents the total amount and 30,000 dollars represents the amount he had the previous year. Is that right?

OpenStudy (malijhaa):

Yes! Correct ! :)

OpenStudy (writerforever18):

Ok, I have that down :)

OpenStudy (writerforever18):

So can I tell you what I think the equation is and you can help me put it together if it's wrong please?

OpenStudy (writerforever18):

What do I have to divide?

OpenStudy (malijhaa):

Oh And Yes Yhu Can Let Me See Yhur Equation

OpenStudy (malijhaa):

I Didnt Mean To Put The Divide Part

OpenStudy (writerforever18):

Oh ok

OpenStudy (malijhaa):

So Whats Yhur Equation?

OpenStudy (writerforever18):

0.09x+0.05y=2060 x+y=30,000

OpenStudy (writerforever18):

Is this correct?

OpenStudy (malijhaa):

Yes Because Yhu Had To Find What The Percentages Represented And Yhu Did. Then Yhu Add Them Together Which Got Yhu To 2060 Then The Variables Equal How Much Money Hans Had :)

OpenStudy (writerforever18):

so do I have to first subtract y on both sides from X+Y=30,000?

OpenStudy (malijhaa):

Yes Correct

OpenStudy (writerforever18):

Ok I am doing this right now and I will tell you what answer I get

OpenStudy (writerforever18):

y=16,000? Is that correct?

OpenStudy (writerforever18):

Which would account 2=16,000

OpenStudy (malijhaa):

Yes!! See, Yhu Dont Need Help :) Yhur Doin Great

OpenStudy (writerforever18):

Haha thank you :) I just never know if I'm getting the problem right, so I always like to make sure. Thank you :) Also, I put 16,000 into the Y variable in "X+Y=30,000" and I got the answer X=14,000. Is that correct?

OpenStudy (malijhaa):

Yup. Thats The Answer! :)

OpenStudy (writerforever18):

I typed in the answer and it was right! Thank you :) Do you think you could please help me with one more?

OpenStudy (writerforever18):

I will post the next question

OpenStudy (malijhaa):

Yeah